OpenStudy (anonymous):

Solve the absolute value equation: |2x-4|=10 A) x=3,-3 B) x = -3, 7 C) x = 3, -7 D) x = 7, -7

OpenStudy (anonymous):

Like with the last one, you need to split it into two separate equations 2x-4 = 10 and the negative version -2x+4 =10 Solving each of those will give you your two values.

OpenStudy (shinalcantara):

taking the absolute value you'll have the equations to find the values of x as: 2x - 4 = 10 and 2x - 4 = -10

OpenStudy (anonymous):

Just for clarity, that works because the absolute value of 2x-4 is the same as the absolute value of -(2x-4) which is -2x+4.

OpenStudy (igreen):

|2x - 4| = 10 Splits into: 2x - 4 = 10 and -2x + 4 = 10 ---------------- 2x - 4 = 10 Add 4 to both sides: 2x = 14 Divide 2 to both sides: x = 7 ---------------- -2x + 4 = 10 Subtract 4 to both sides: -2x = 6 Divide -2 to both sides: What's 6 / -2?

OpenStudy (igreen):

@pigglet

OpenStudy (anonymous):

-3, sorry, got distracted.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, so our answers are 7 and -3.

OpenStudy (anonymous):

so, B. thanks!